# Quantum correction to the orbital Hall effect

### v1

@@ Abstract
The orbital Hall effect (OHE) typically is evaluated using an intuitive approach based on semiclassical considerations. Here we provide a full quantum mechanical evaluation of the orbital current, including all matrix elements of the position operator. We recover previous results and find an additional quantum correction due to the non-commutativity of the position and velocity operators. We evaluate the full OHE for the topological antiferromagnet CuMnAs showing that the quantum correction dominates the response and changes its sign. We discuss implications for experiment.
@@
Introduction. The non-equilibrium properties of Bloch electrons’ orbital angular momentum (OAM) Yafet (1963); Vanderbilt (2018) have come under intense scrutiny with the advent of orbitronics Bernevig et al. (2005); Das (2023); Rhonald Burgos Atencia and Culcer (2024), whose focus is generating non-equilibrium OAM densities and currents Choi et al. (2023); Liu and Culcer (2024); Xiao et al. (2021); Pezo et al. (2022); Salemi and Oppeneer (2022); Cysne et al. (2021); Luis M. Canonico (2024); Go et al. (2018); Jo et al. (2018); Sala and Gambardella (2022a); Wang et al. (2023a); Lyalin et al. (2023); Dutta and Tulapurkar (2022); Sala et al. (2023); Zhang et al. (2022); Ding et al. (2022); Voss et al. (2024). A significant technological motivation behind recent efforts is the electrical manipulation of magnetic degrees of freedom Kim et al. (2021); Chen et al. (2018); Go and Lee (2020); Lee et al. (2021a); Zheng et al. (2020); Lee et al. (2021b, c); Ding et al. (2020); Sala and Gambardella (2022b); Hayashi et al. (2023); Li et al. (2023), with an emphasis on weakly spin-orbit coupled materials Salvador-Sánchez et al. (2024); Seifert et al. (2023); Ünzelmann et al. (2020); Salemi et al. (2019); Tang and Bauer (2024); Wang et al. (2023b); Busch et al. (2024). From a fundamental science perspective, the OHE has been proposed as one of the mechanisms behind the observed valley Hall effect Bhowal and Vignale (2021); Wu et al. (2013); Song et al. (2015); Wu et al. (2013); Bhowal and Satpathy (2020); Xue et al. (2020), and attempts to disentangle these two effects are now underway Cysne et al. (2021, 2022); Kazantsev et al. (2024).
@@
Determining the OAM and OHE involves working with the position operator, which is challenging in extended systems Fonseca et al. (2023); Ado et al. (2024); Yoda et al. (2018, 2015); Park et al. (2012, 2013, 2011). The main obstacle is the band-diagonal elements of the position operator. This obstacle can be circumvented in equilibrium, since the equilibrium OAM involves only inter-band matrix elements of the position operator. At present, the approach universally adopted for non-equilibrium systems is to start from the equilibrium OAM matrix elements and combine those with a non-equilibrium distribution found using standard methods such as the Boltzmann equation or the Kubo formula Rhonald Burgos Atencia and Culcer (2024); Liu and Culcer (2024); Cysne et al. (2022); Bhowal and Vignale (2021); Cysne et al. (2021); Pezo et al. (2022, 2023); Hanke et al. (2016); Go et al. (2020, 2018); Jo et al. (2018); Baek and Lee (2021); Kim et al. (2021); Tang and Bauer (2024); Inoue et al. (2004, 2006); Chen et al. (2024); Costa et al. (2023); Barbosa et al. (2023). This includes earlier work by some of us Liu and Culcer (2024). We argue that this approach, which has never been justified on fundamental grounds, is incomplete and misses important terms, a fact that can be seen already at the operator level. In the modern theory of orbital magnetization Vanderbilt (2018) the OAM is expressed as the expectation value of the symmetrized vector product of the position and velocity operators. To be consistent with the modern theory, the orbital current needs to be evaluated as the full quantum mechanical expectation value of the product of the OAM and velocity operators, including the band-diagonal terms in the position operator, and this needs to be found explicitly for a non-equilibrium system.
@@
Figure 1: The intrinsic OHE for CuMnAs [Eq. 6] with OAM polarization along $\hat{\bm{z}}$ and electric field along $\hat{\bm{x}}$. The orbital current $j^{\vphantom{\dagger}}_{L}$ is given by the conventional contribution $j_{\rm conv}$ plus the quantum correction $\Delta j$. The parameters are $t=0.08\,{\rm eV}$ and $\tilde{t}=1\,{\rm eV}$, $\alpha_{\rm R}=0.8,\alpha_{\rm D}=0$,and ${\bm{h}}_{\rm AFM}=(0.85,0,0)\,{\rm eV}$.
@@
In this paper we present a full quantum mechanical evaluation of the non-equilibrium expectation value of the orbital current operator, including all matrix elements of the position operator. Our principal finding is the full expression for the orbital current in response to an electric field, which is given by $j^{\vphantom{\dagger}}_{L}=j_{\rm conv}+\Delta j$ and is shown in Fig. 1 for CuMnAs. Here $j_{\rm conv}$ is the conventional orbital current, which has been evaluated to date, and is based on the equilibrium OAM and the non-equilibrium distribution. The quantum correction $\Delta j$ consists of contributions from: (i) the electron group velocity leading to band off-diagonal matrix elements of the OAM; (ii) inter-band coherence terms induced by the band-diagonal part of the position operator; and (iii) the non-commutativity of the position and velocity operators. We develop a quantum kinetic equation for the density matrix and evaluate the full OHE for the tight-binding band structure of the topological antiferromagnet CuMnAs, where $\Delta j$ provides the dominant contribution to the OHE, such that the resulting OHE has the opposite sign to the conventional formula. Our result is general and applies to all solids, having important implications for the interpretation of experiments. We focus on the intrinsic case and in closing discuss generalizations to disordered systems under certain approximations. We stress that our concern is with the evaluation of the orbital current, not its definition, which we take to be the usual product of the OAM and velocity operators.
@@
OAM and orbital current operators. We begin by considering a generalized dipole operator $\hat{\bm{d}}^{A}={\textstyle{1\over 2}}\{\hat{A},\hat{\bm{r}}\}$, where $\hat{A}$ represents a general operator which may be a scalar such as charge, a vector component of spin, velocity, etc. The expectation value of such an operator is Tr$(\hat{\bm{d}}^{A}\hat{\rho})$ with $\hat{\rho}$ the density matrix. We define the OAM operator as the symmetrized combination $\hat{\bm{L}}={\textstyle{1\over 2}}(\hat{\bm{r}}\times\hat{\bm{v}}+\hat{\bm{v}}\times\hat{\bm{r}})$, with $\hat{\bm{v}}$ the velocity operator, and we work in the Hilbert space spanned by Bloch wave-functions ${\big{|}\hskip 1.00006pt\Psi_{m{\bm{k}}}\hskip 1.00006pt\big{\rangle}}=e^{i{\bm{k}}\cdot{\bm{r}}}{\big{|}\hskip 1.00006ptu_{m{\bm{k}}}\hskip 1.00006pt\big{\rangle}}$. Determining the OAM expectation value is tantamount to substituting the appropriate velocity component for $\hat{A}$ above, whereupon the correct expression for the equilibrium OAM in the modern theory immediately emerges.
@@
where ${\mathcal{R}}_{\beta}^{lm}=i\,{\langle\hskip 1.00006ptu_{l}\hskip 1.00006pt|\hskip 1.00006pt{{\partial}u_{m}\over{\partial}k_{\beta}}\hskip 1.00006pt\rangle}$ is the Berry connection, which reflects the wave vector dependence of the basis functions. The velocity operator in turn has matrix elements
$\hbar\,[{\hat{v}}_{\delta}]^{mn}_{\bm{k}}={{\partial}{\varepsilon}_{m}\over{\partial}k_{\delta}}\,\delta_{mn}+i\,{\mathcal{R}}^{mn}_{\delta}({\varepsilon}_{m}-{\varepsilon}_{n})\ ,$ (3)
where ${\varepsilon}_{m}={\varepsilon}_{m{\bm{k}}}$ is the band energy at wavevector ${\bm{k}}$.
To evaluate the full OHE we require the non-equilibrium correction to the density matrix in an electric field, for which we turn to linear response theory following the approach of Refs. Culcer et al. (2017); Atencia et al. (2022). The single-particle density operator obeys the quantum Liouville equation, ${\partial}{\hat{\rho}}/{\partial}t+(i/\hbar)[{\hat{H}},{\hat{\rho}}]=0$, where ${\hat{H}}={\hat{H}}_{0}+{\hat{H}}_{E}$. Here ${\hat{H}}_{0}$ is the band Hamiltonian and ${\hat{H}}_{E}=e{\bm{E}}\cdot\hat{\bm{r}}$ is potential due to the external electrical field. At this stage we focus on intrinsic effects and do not consider disorder scattering, which will be discussed in closing. In the crystal momentum representation the equilibrium density matrix has the diagonal form $\rho_{0{\bm{k}}}^{mn}=f_{m}\,\delta_{mn}$, where $f_{m}\equiv f({\varepsilon}_{m{\bm{k}}})$ is the Fermi-Dirac distribution for band $m$. In an electric field the density matrix can be written as $\hat{\rho}=\rho_{0}+\rho_{E}$, and, in linear response, it has been shown that Culcer et al. (2017)
@@
Figure 2: The quantum correction $\Delta j$ is decomposed as $\Delta j=\Delta{j_{1}}+\Delta{j_{2}}+\Delta{j_{3}}$. $\Delta j_{1}$ is from the group velocity, $\Delta j_{2}$ is from the band-diagonal part of position operator, and $\Delta j_{3}$ is from the non-commutativity of the position and velocity operators. We have used the same parameters as in Fig. 1.
Once $\rho^{mn}_{E{\bm{k}}}$ is found the quantum correction can be written as $\Delta j=\Delta j_{1}+\Delta j_{2}+\Delta j_{3}$, where
$\Delta j^{\alpha}_{\delta,1}$ $={\textstyle{1\over 2}}\,\epsilon_{\alpha\beta\gamma}\sum_{m,{\bm{k}}}\big{\{}{\mathcal{R}}^{\vphantom{\dagger}}_{\beta},\rho^{\vphantom{\dagger}}_{E{\bm{k}}}\big{\}}^{mm}\,v^{mm}_{\delta}v^{mm}_{l}$ (5)
$\Delta j^{\alpha}_{\delta,2}$ $={\textstyle{i\over 4}}\,\epsilon^{\vphantom{*}}_{\alpha\beta\gamma}\!\sum_{m,n,{\bm{k}}}{2eE_{\mu}\left[{D\Xi^{0}_{\beta}\over Dk_{\mu}}\right]^{mn}\!\!\!+\big{\{}\hbar v^{\vphantom{\dagger}}_{\beta},\rho^{\vphantom{\dagger}}_{E{\bm{k}}}\big{\}}^{mn}\over{\varepsilon}_{n}-{\varepsilon}_{m}}\{v_{\gamma},v_{\delta}\}^{nm}$
@@
where we have abbreviated $\big{[}\Xi^{0}_{\beta}\big{]}^{mn}={\textstyle{1\over 2}}{\mathcal{R}}^{mn}_{\beta}(f_{m}+f_{n})$, and $m\neq n$ is understood in all the summations. The result is gauge invariant, as shown in the Supplement. The OAM polarization is taken to be along the $\alpha$-direction while the transport direction is denoted by $\delta$. The three terms entering the quantum correction reflect inter-band coherence induced by the electric field. $\Delta j_{1}$ has the same structure as $j_{\rm conv}$, except that the electron group velocity appears instead of the inter-band velocity. $\Delta j_{2}$ is the result of restoring the band-diagonal matrix elements of the position operator on the left side of Eq. 1. Finally, $\Delta j_{3}$ reflects the non-commutativity of the position and velocity operators, which will be discussed below.
Tetragonal CuMnAs. We now exemplify our evaluation by studying the OHE in a tight-binding model of CuMnAs. CuMnAs can exist in both orthorhombic and tetragonal crystal structures. Both have been identified experimentally Linn et al. (2023); Wadley et al. (2015), and both can host the antiferromagnetic Dirac semimetal phase. Here we take the stable tetragonal CuMnAs phase as an example, which has opposite spins lying on a bipartite lattice. Such an arrangement breaks the $\mathcal{P}$ as well as the $\mathcal{T}$ symmetry locally. However, the combined $\mathcal{PT}$ symmetry is preserved by the exchange of the sublattices with the flip of oppositely aligned spins Das et al. (2023); Watanabe and Yanase (2021a), enforcing the bands at each momentum is doubly degenerate Li et al. (2020). The tight-binding Hamiltonian for CuMnAs without external magnetization is written as Wang et al. (2021)
$H_{0{\bm{k}}}=\begin{pmatrix}{\varepsilon}_{0}({\bm{k}})+{\bm{h}}({\bm{k}})\cdot{\bm{\sigma}}&V_{\rm AB}({\bm{k}})\\ V_{\rm AB}({\bm{k}})&{\varepsilon}_{0}({\bm{k}})-{\bm{h}}({\bm{k}})\cdot{\bm{\sigma}}\end{pmatrix}\quad,$ (6)
Here ${\varepsilon}_{0}({\bm{k}})=-t(\cos k_{x}+\cos k_{y})$ and $V_{\text{AB}}=-2\tilde{t}\cos(k_{x}/2)\cos(k_{y}/2)$, where $t$ and $\tilde{t}$ denote hopping between orbitals of the same and different sublattices, respectively. The sublattice-dependent spin-orbit coupling and the magnetization field are included in
@@
with $\alpha^{\vphantom{\dagger}}_{\rm R}$ and $\alpha^{\vphantom{\dagger}}_{\rm D}$ the Rashba and Dresselhaus type spin-orbit coupling coefficients Watanabe and Yanase (2021b). In the band dispersion gapless points appear along the high symmetry line $(k_{x}=\pi)$, and we consider the local topology around the high-symmetry point. We expand the Hamiltonian in the vicinity of $(k_{x},k_{y})=(1,0.5)\pi$ up to $\delta k_{x,y}=\pm 0.1\pi$. Then we numerically integrate the full OHE for the approximated model Hamiltonian around $(1,0.5)\pi$ up to ±$\pm 0.1\pi$ by ${\bm{k}}$-discretization. The intrinsic anomalous Hall effect is proportional to the integration over the Fermi sea of the Berry curvature of each occupied band or equivalently to the integral of the Berry phases over cuts of the Fermi-surface segments, which ensures Fermi sea effects are included in the Fermi-liquid picture in order to make the integration tractable Haldane (2004); Wang et al. (2007). Here the Fermi surface of CuMnAs model is not isotropic. We interpret the full OHE as a Fermi sea effect. The transformation of the Fermi sea to Fermi surface calculation does not simplify the full OHE integral. We calculate the orbital Hall conductivity dependence on the Fermi energy $E_{\rm F}$, when $E_{\rm F}$ is in the conduction band. For the tight-binding Hamiltonian of CuMnAs, we obtain the total intrinsic OHE in Fig. 1, showing that the magnitude of both contributions decreases with increasing $E_{\rm F}$ Liu and Culcer (2024); Bhowal and Vignale (2021). Remarkably, the conventional contribution corresponds to only one fraction of the total orbital current, and the quantum correction $\Delta j$ accounts for almost the entire value. We also plot the individual contributions to $\Delta j$ in Fig. 2, showing that in this model the dominant term is $\Delta j_{3}$, which arises from the non-commutativity of the position and velocity operators.
Discussion. We begin with the rationale behind our approach. According to the philosophy of quantum mechanics, the starting point in any calculation is the operator under consideration. This approach has underpinned the entire modern theory of orbital magnetization, which takes the OAM operator as the starting point in the evaluation of the OAM. The equilibrium OAM stems from the inter-band matrix elements of the position and velocity operators, in fact it can be regarded as an inter-band coherence phenomenon in the sense of Ref. Culcer et al. (2017), albeit an equilibrium version of it. Traditionally these inter-band matrix elements were neglected in transport calculations Ashcroft and Mermin (1976); Cardona (1969), and the OAM would have continued to be missed if this neglect persisted. According to this reasoning, the starting point of the orbital current calculation is the orbital current operator, whose matrix elements must be evaluated in their entirety. Such a full quantum mechanical evaluation reveals the presence of several contributions to the orbital current, which affect the composition and magnitude of the orbital Hall effect. For the model system chosen the quantum correction is much larger in magnitude than the conventional value. In CuMnAs the conventional and corrected OHE also differ by a sign, yet there is no guarantee that such a difference exists in general. To determine the sign in principle the OHE can be imaged using the techniques developed by Zeldov, although this remains to be demonstrated Grover et al. (2022).
The equilibrium OAM is expressed purely in terms of off-diagonal matrix elements of the position and velocity operators Sundaram and Niu (1999); Chang and Niu (1995, 1996), hence the unruly position operator is tamed Thonhauser et al. (2005); Ceresoli et al. (2006); Resta (2020). Nevertheless, as our Supplement shows, the presence of only inter-band position and velocity matrix elements in the equilibrium OAM is a result of the equilibrium density matrix being purely diagonal in the band index. In a non-equilibrium scenario the evaluation of the OAM and related quantities such as the orbital current may contain band-diagonal matrix elements of the position and velocity operators.

### v2

# Quantum corrections to the orbital Hall effect

@@ Abstract
Evaluations of the orbital Hall effect (OHE) have only retained inter-band matrix elements of the position operator. Here we evaluate the OHE including all matrix elements of the position operator, including the technically challenging intra-band elements. We recover previous results and find quantum corrections due to the non-commutativity of the position and velocity operators and inter-band matrix elements of the orbital angular momentum. The quantum corrections dominate the OHE responses of the topological antiferromagnet CuMnAs and of massive Dirac fermions.
@@
Introduction. The non-equilibrium properties of Bloch electrons’ orbital angular momentum (OAM) [1, 2] have come under intense scrutiny with the advent of orbitronics [3, 4, 5], whose focus is generating non-equilibrium OAM densities and currents [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. The orbital Hall effect (OHE) refers to the generation of a transverse OAM current by an electric field. The technological motivation underpinning OHE efforts is the electrical manipulation of magnetic degrees of freedom [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33], with an emphasis on weakly spin-orbit coupled materials [34, 35, 36, 37, 38, 39, 40]. From a fundamental science perspective, the OHE has also been proposed as one of the mechanisms behind the observed valley Hall effect [41, 42, 43, 42, 44, 45], and attempts to disentangle these two effects are now underway [11, 46, 47].
@@
The OAM and OHE involve the position operator, which is challenging in extended systems [48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58]. In particular, the band-diagonal elements of the position operator are themselves differential operators, which act on the quantities surrounding them. This difficulty can be circumvented in equilibrium, since the equilibrium OAM expectation value involves only inter-band matrix elements of the position operator. In light of this, the conventional evaluation of the OHE, which includes earlier work by some of us [7], has proceeded as follows: (i) start with the OAM matrix elements appearing in the equilibrium OAM expectation value; (ii) multiply these by the velocity operator matrix elements; and (iii) combine this product with a non-equilibrium distribution found using standard methods such as the Boltzmann equation or the Kubo formula [5, 7, 46, 41, 11, 9, 59, 60, 61, 13, 14, 62, 23, 38, 63, 64, 65, 66, 67]. We argue that this approach is incomplete and misses important terms, a fact that can be seen already at the operator level. Replacing the OAM operator appearing inside the orbital current by the equilibrium OAM matrix elements is inappropriate, as it amounts to neglecting the band-diagonal matrix elements of the position operator. These matrix elements, being differential operators, act on the velocity matrix elements surrounding them. The additional resulting terms are substantial, and in certain cases overwhelm the conventional terms in the OHE.
@@
Figure 1: The intrinsic OHE for CuMnAs [Eq. 6] with OAM polarization along $\hat{\bm{z}}$ and electric field along $\hat{\bm{x}}$. The inset plots the conventional contribution $j_{\rm conv}$ again, with a different scale on the y-axis. The parameters are $t=0.08\,{\rm eV}$ and $\tilde{t}=1\,{\rm eV}$, $\alpha_{\rm R}=0.8,\alpha_{\rm D}=0$,and ${\bm{h}}_{\rm AFM}=(0.85,0,0)\,{\rm eV}$.
@@
In this paper we present a full quantum mechanical evaluation of non-equilibrium orbital current accounting for all matrix elements of the position operator, including its intra-band matrix elements, which are technically challenging. Our principal finding is the full expression for the orbital current in response to an electric field, given by $j^{\vphantom{\dagger}}_{L}=j_{\rm conv}+\Delta j$. Here $j_{\rm conv}$ is the conventional orbital current, which has been evaluated to date, and is based on the equilibrium OAM and the non-equilibrium distribution. The quantum corrections $\Delta j$ consist of contributions from: (i) the electron group velocity leading to band off-diagonal matrix elements of the OAM; (ii) inter-band coherence terms induced by the band-diagonal part of the position operator; and (iii) the non-commutativity of the position and velocity operators. We develop a quantum kinetic equation for the density matrix and determine a general expression for the OHE applicable to an arbitrary band structure, focussing on the intrinsic case. Our result, presented in Eqs. 4-5 below, is general and applies to all solids, having important implications for the interpretation of experiments. To illustrate this, we first evaluate the full OHE for the topological antiferromagnet CuMnAs, shown in Fig. 1. It is seen that $\Delta j$ provides the dominant contribution to the OHE, and the resulting OHE has the opposite sign to the conventional formula. We also evaluate the full OHE for a model of massive Dirac fermions, where again the quantum corrections dominate the response, and are also influenced by tilting and warping effects. We stress that our concern is with the evaluation of the orbital current, not its definition, which we take to be the usual product of the OAM and velocity operators.
@@
OAM and orbital current operators. We begin by considering a generalized dipole operator $\hat{\bm{d}}^{A}={\textstyle{1\over 2}}\{\hat{A},\hat{\bm{r}}\}$, where $\hat{A}$ represents a general operator. The expectation value of such an operator is Tr$(\hat{\bm{d}}^{A}\hat{\rho})$ with $\hat{\rho}$ the density matrix. We define the OAM operator as the symmetrized combination $\hat{\bm{L}}={\textstyle{1\over 2}}(\hat{\bm{r}}\times\hat{\bm{v}}-\hat{\bm{v}}\times\hat{\bm{r}})$, with $\hat{\bm{v}}$ the velocity operator, and we work in the Hilbert space spanned by Bloch wave-functions ${\big{|}\hskip 1.00006pt\Psi_{m{\bm{k}}}\hskip 1.00006pt\big{\rangle}}=e^{i{\bm{k}}\cdot{\bm{r}}}{\big{|}\hskip 1.00006ptu_{m{\bm{k}}}\hskip 1.00006pt\big{\rangle}}$. Determining the OAM expectation value is tantamount to substituting the appropriate velocity component for $\hat{A}$ above, whereupon the correct expression for the equilibrium OAM in the modern theory immediately emerges.
@@
where ${\mathcal{R}}_{\beta}^{lm}=i\,{\langle\hskip 1.00006ptu_{l}\hskip 1.00006pt|\hskip 1.00006pt{{\partial}u_{m}\over{\partial}k_{\beta}}\hskip 1.00006pt\rangle}$ is the Berry connection, which reflects the wave vector dependence of the basis functions. The velocity operator in turn has matrix elements $\hbar\,[{\hat{v}}_{\delta}]^{mn}_{\bm{k}}=\partial{\varepsilon}_{m}/\partial k_{\delta}\,\delta_{mn}+i\,{\mathcal{R}}^{mn}_{\delta}({\varepsilon}_{m}-{\varepsilon}_{n})$, where ${\varepsilon}_{m}={\varepsilon}_{m{\bm{k}}}$ is the band energy at wavevector ${\bm{k}}$.
Orbital Hall Effect: General expression. To evaluate the OHE we require the non-equilibrium correction to the density matrix in an electric field, for which we turn to linear response theory following the approach of Refs. [68, 69]. The single-particle density operator obeys the quantum Liouville equation, ${\partial}{\hat{\rho}}/{\partial}t+(i/\hbar)[{\hat{H}},{\hat{\rho}}]=0$, where ${\hat{H}}={\hat{H}}_{0}+{\hat{H}}_{E}$. Here ${\hat{H}}_{0}$ is the band Hamiltonian and ${\hat{H}}_{E}=e{\bm{E}}\cdot\hat{\bm{r}}$ is potential due to the external electrical field. At this stage we focus on intrinsic effects and do not consider disorder scattering. In the crystal momentum representation the equilibrium density matrix has the diagonal form $\rho_{0{\bm{k}}}^{mn}=f_{m}\,\delta_{mn}$, where $f_{m}\equiv f({\varepsilon}_{m{\bm{k}}})$ is the Fermi-Dirac distribution for band $m$. In an electric field the density matrix can be written as $\hat{\rho}=\rho_{0}+\rho_{E}$, and, in linear response, it has been shown that [68]
@@
Figure 2: The decomposed quantum correction $\Delta j$ for CuMnAs with OAM polarization along $\hat{\bm{z}}$ and electric field along $\hat{\bm{x}}$. We have used the same parameters as in Fig. 1.
Once $\rho^{mn}_{E{\bm{k}}}$ is found the conventional orbital current is
$j_{\rm conv}\displaystyle={\textstyle{1\over 4}}\epsilon_{\alpha\beta\gamma}\sum_{m,{\bm{k}}}\big{\{}{\mathcal{R}}^{\vphantom{\dagger}}_{\beta},\rho^{\vphantom{\dagger}}_{E{\bm{k}}}\big{\}}^{mm}\,\big{\{}v_{\delta},v_{\gamma}\big{\}}^{mm}\,,$ (4)
where only band off-diagonal components of the velocity operators enter. The quantum corrections can be written as $\Delta j=\Delta j_{1}+\Delta j_{2}+\Delta j_{3}$, where
$\Delta j^{\alpha}_{\delta,1}$ $={\textstyle{1\over 2}}\epsilon_{\alpha\beta\gamma}\sum_{m,{\bm{k}}}\big{\{}{\mathcal{R}}^{\vphantom{\dagger}}_{\beta},\rho^{\vphantom{\dagger}}_{E{\bm{k}}}\big{\}}^{mm}\,v^{mm}_{\delta}v^{mm}_{\gamma}$ (5)
$\Delta j^{\alpha}_{\delta,2}$ $={\textstyle{i\over 4}}\epsilon_{\alpha\beta\gamma}\!\!\!\sum_{m,n,{\bm{k}}}\!\!\!{2eE_{\mu}\left[{D\Xi^{0}_{\beta}\over Dk_{\mu}}\right]^{mn}\!\!\!\!+\big{\{}\hbar v^{\vphantom{\dagger}}_{\beta},\rho^{\vphantom{\dagger}}_{E{\bm{k}}}\big{\}}^{mn}\over{\varepsilon}_{n}-{\varepsilon}_{m}}\{v_{\gamma},v_{\delta}\}^{nm}$
@@
where we have abbreviated $\big{[}\Xi^{0}_{\beta}\big{]}^{mn}={\textstyle{1\over 2}}{\mathcal{R}}^{mn}_{\beta}(f_{m}+f_{n})$, and $m\neq n$ is understood in all the summations. These expressions are general and apply to arbitrary band structures, provided the velocity and Berry connection matrix elements can be calculated. The result is gauge invariant, as shown in the Supplement[70]. The OAM polarization is taken to be along the $\alpha$-direction while the transport direction is denoted by $\delta$. We stress that $\Delta j$ has never been reported previously. The three terms entering the quantum correction $\Delta j$ reflect inter-band coherence induced by the electric field. $\Delta j_{1}$ has the same structure as $j_{\rm conv}$, except that the electron group velocity appears instead of the inter-band velocity. $\Delta j_{2}$ is the result of restoring the band-diagonal matrix elements of the position operator on the left side of Eq. 1. Finally, $\Delta j_{3}$ reflects the non-commutativity of the position and velocity operators.
Tetragonal CuMnAs. We now exemplify our evaluation by studying the OHE in a tight-binding model of CuMnAs. CuMnAs can exist in both orthorhombic and tetragonal crystal structures [77, 78], both can host the antiferromagnetic Dirac semimetal phase. Here we take the stable tetragonal CuMnAs phase as an example, which has opposite spins lying on a bipartite lattice. Such an arrangement preserves the combined $\mathcal{PT}$ symmetry by the exchange of the sublattices with the flip of oppositely aligned spins [79, 80], enforcing that bands at each momentum are doubly degenerate [81]. The tight-binding Hamiltonian for CuMnAs without external magnetization is written as [82]
$H_{0,{\bm{k}}}=\begin{pmatrix}{\varepsilon}_{0}({\bm{k}})+{\bm{h}}({\bm{k}})\cdot{\bm{\sigma}}&V_{\rm AB}({\bm{k}})\\ V_{\rm AB}({\bm{k}})&{\varepsilon}_{0}({\bm{k}})-{\bm{h}}({\bm{k}})\cdot{\bm{\sigma}}\end{pmatrix}\quad,$ (6)
where ${\varepsilon}_{0}({\bm{k}})=-t(\cos k_{x}+\cos k_{y})$ and $V_{\text{AB}}=-2\tilde{t}\cos(k_{x}/2)\cos(k_{y}/2)$, where $t$ and $\tilde{t}$ denote hopping between orbitals of the same and different sublattices, respectively. The sublattice-dependent spin-orbit coupling and the magnetization field are included in
@@
with $\alpha^{\vphantom{\dagger}}_{\rm R}$ and $\alpha^{\vphantom{\dagger}}_{\rm D}$ the Rashba and Dresselhaus type spin-orbit coupling coefficients [83]. In the band dispersion gapless points appear along the high symmetry line $(k_{x}=\pi)$. We expand the Hamiltonian in the vicinity of $(k_{x},k_{y})=(1,0.5)\pi$ up to $\delta k_{x,y}=\pm 0.1\pi$. Then we numerically integrate the full OHE for the approximated model Hamiltonian around $(1,0.5)\pi$ up to ±$\pm 0.1\pi$ by ${\bm{k}}$-discretization. This is justified since the OAM itself is strongly peaked around these points. We calculate the orbital Hall conductivity dependence on the Fermi energy $E_{\rm F}$, when $E_{\rm F}$ is in the conduction band, $E_{\rm F}=0$ corresponds to the middle of the band gap. For CuMnAs we obtain the total intrinsic OHE in Fig. 1.  [^1] Note: our plotted values of the orbital conductivity have been multiplied by $-1$, this is for the added factor of the electron charge in the orbital magnetic moment, such that our results are comparable with Ref. [41].. Remarkably, the conventional contribution corresponds to only one fraction of the total orbital current, while the quantum correction $\Delta j$ accounts for almost the entire value, which has the opposite sign. To determine the sign the OHE could in principle be imaged using the techniques developed for OAM densities [85]. The current is non-zero in the gap due to the Fermi sea contribution from the filled valence band states. We also plot the individual contributions to $\Delta j$ in Fig. 2, showing that in this model the dominant term is $\Delta j_{3}$, which arises from the non-commutativity of the position and velocity operators.
Dirac Model. We look to further emphasise the importance of the quantum correction in the orbital current by examining a simple and general two-band model – a massive tilted Dirac cone with trigonal warping
$H_{0,\bm{k}}=$ $\alpha(k_{y}\sigma_{x}-k_{x}\sigma_{y})+\kappa((k_{x}^{2}-k_{y}^{2})\sigma_{x}-2k_{x}k_{y}\sigma_{y})$ (8)
$+m\sigma_{z}+tk_{x}\sigma_{0}\,,$
where $\kappa$ is the warping parameter, $t$ is the tilting parameter and, $\sigma_{x,y,z,0}$ are the Pauli matrices. If one first ignores the warping and tilt terms, then for the Fermi energy $E_{\rm F}$ in the conduction band the orbital current is
$\begin{split}j_{{\rm conv},y}^{z}&={eE_{x}m^{2}\alpha^{2}\over 24\pi\hbar^{2}E_{\rm F}^{3}}\\ \Delta j_{1,y}^{z}&={eE_{x}\alpha^{2}(3E_{\rm F}^{2}-m^{2})\over 48\pi\hbar^{2}E_{\rm F}^{3}}\quad.\end{split}$ (9)
In this simple case $\Delta j_{2,y}^{z}$ and $\Delta j_{3,y}^{z}$ are zero. At the conduction band bottom $E_{F}\rightarrow m$ and the two contributions are equal. However, as the Fermi energy increases the conventional contribution decays rapidly, whereas the quantum correction decays much more slowly and dominates. When $E_{\rm F}$ is in the gap the orbital current will be quantized, the total orbital current will be $j_{L}=eE_{x}\alpha^{2}/12\pi\hbar^{2}m$. The same behaviors are also seen in Fig.3 when including both the warping and tilt. When warping and the tilt are included we find $\Delta j_{2,y}^{z}$ and $\Delta j_{3,y}^{z}$ to be non-zero, yet they remain relatively small. Here $\Delta j_{1,y}^{z}$ is the dominant contribution to the orbital current. Hence, even for this simple model we find that the conventional method for the evaluation of the orbital current is insufficient and the quantum correction is required for an accurate calculation.
Figure 3: The intrinsic $j_{L}$ and each of its components for our two-band Dirac model. Here we use $\alpha=4$ eVÅ, $m=0.01$ eV, $t=0.2\alpha$ and $\kappa=0.8$ eVÅ2.
Discussion. We have shown that the non-equilibrium expectation value of the orbital current operator contains terms that have not been evaluated previously, which, for the models considered, have substantial magnitudes. How does our approach differ from what has been done to date? To answer this, we note that the starting point of any quantum mechanical calculation is the operator under consideration. The OAM itself is the product of the position and velocity operators. The equilibrium OAM stems from the inter-band matrix elements of these operators, which were traditionally neglected [86, 87], and the OAM would have continued to be missed if this neglect persisted. According to the same reasoning, the starting point of any calculation of the OHE must be the orbital current operator, that is, the product of the OAM and velocity operators. The non-equilibrium expectation value of the orbital current operator will contain intra-band and inter-band matrix elements of both the position and velocity operators. To date, only the inter-band position matrix elements were retained in the OHE, while the intra-band position matrix elements were neglected.
The reason for this neglect is straightforward. Quantum mechanically, in principle, once the matrix elements are known the operator is known. It is natural to assume that constructing the orbital current operator matrix elements involves multiplying the known OAM matrix elements with the known velocity matrix elements, and extracting the OAM matrix elements from the equilibrium expectation value of the OAM. This procedure, however, is problematic, because of two reasons. Firstly, the matrix elements of the position operator for Bloch electrons are differential operators. Secondly, the equilibrium density matrix only selects one part of the OAM matrix elements. The equilibrium OAM expectation value involves only off-diagonal matrix elements of the position and velocity operators [88, 89, 90]. The net result is that intra-band matrix elements of the position operator are thrown away when the OAM expectation value is taken in equilibrium. Hence, the OAM matrix elements extracted from the equilibrium OAM expectation value are not sufficient to determine the orbital current. The intra-band matrix elements of the position operator must be included, and we have developed a powerful technique for handling them. We have focussed on the Bloch picture of electrons in crystals. Different insights into the OAM emerge from the Wannier and Bloch formulations [5, 46, 9, 59, 91, 92]. Since an exact correspondence exists between the Bloch and Wannier bases, the fundamental considerations outlined in this work will be applicable to Wannier studies as well.